# Simultaneously Approximating All Norms for Massively Parallel Correlation Clustering

### v1

@@ 1 Introduction
In light of the networks growth, it is imperative to develop efficient parallel algorithms. This urgency is particularly pronounced in machine learning and data mining applications, where timely and efficient processing is essential for extracting meaningful insights from vast datasets. Many works in the literature aim to design efficient parallel algorithms [BFS12, CDK14, PPO+15, FN19, CCMU21, CALM+21, AW22, CKL+24, CHS24]. The MPC model, as a theoretical abstraction of several real-world parallel models such as MapReduce [DG08], is a prevalent methodology employed in these works.
@@ 1.2 Overview of Techniques
By [CS19], we can reduce the problem of approximating all monotone symmetric norms to approximating all top-$k$ norms. We then construct a fractional solution $x$, which is a metric over $V$ with range $[0,1]$, such that the fractional disagreement vector for $x$ has top-$k$ norm at most $12.66\cdot{\mathrm{opt}}_{k}$ for any $k\in[n]$, where ${\mathrm{opt}}_{k}$ is the cost of the optimum clustering under the top-$k$ norm. Then, we can use the 5-approximate rounding algorithm of KMZ [KMZ19], to obtain a simultaneous $63.3$-approximation for all top-$k$ norms. The KMZ rounding algorithm have two crucial properties that we need: it does not depend on $k$ and it achieves a per-vertex guarantee.
We elaborate more on how to construct the metric $x:{V\choose 2}\to[0,1]$. A natural idea to assign the LP values, that was used by [DMN24], is to set $x_{uv}$ based on the intersection of the neighborhood between $u$ and $v$. Intuitively, the more common neighbors two nodes share, the closer they should be. A straightforward approach to implement this idea is to set $x_{uv}=1-\frac{|N(u)\cap N(v)|}{\max(d(u),d(v))}$, where $N(u)$ denotes the neighboring nodes of $u$ in $G$ and $d(u)=|N(u)|$ denotes the degree of $u$; it is convenient to assume $u\in N(u)$. This approach works for the top-$1$ norm (i.e., the $\ell_{\infty}$ norm) as discussed in [DMN23], but fails for the top-$n$ norm (i.e., the $\ell_{1}$ norm). Consider a star graph, where the optimal clustering under the top-$n$ norm has a cost of $n-2$. This approach will assign $x_{uv}=1-\frac{1}{2}=1/2$ for all --edges, leading to an LP cost of $\Theta(n^{2})$ and a gap of $\Omega(n)$. [DMN24] addressed the issue by rounding up LP values to $1$ for --edge, if for a given node, its total --edges LP disagreement is larger than the number of its $+$edges. After the transformation, the triangle inequalities are only satisfied approximately, but this can be handled with $O(1)$ loss in the approximation ratio.
@@
The challenge to implement the algorithm in MPC model is the sequential nature of the algorithm. [KMZ19] observe that in each round, if we select the nodes that maximize $L(u)=\sum_{x_{uv}\leq r}(r-x_{uv})$ as cluster center, we can effectively bound each node’s algorithmic cost, where $r$ is the final ratio. However, choosing a node that maximizes some target inherently makes the process sequential. Our key observation is that, instead of selecting the node that maximizes $L(u)$, we can allow some approximation. This strategy still permits achieving a reasonable approximate ratio with an additional $1+\epsilon$ overhead while allowing the selection of multiple nodes as cluster centers, thereby parallelizing the rounding process. In each round, there might be several candidate cluster centers with conflicts. To resolve these conflicts, we employ the classical Luby’s algorithm [Lub85, CDK14] to find a maximal independent set, ensuring that none of the cluster centers have conflicts.
@@
We give some preliminaries in Section 2. In Section 3, we describe our simultaneous $O(1)$-approximation algorithm for correlation clustering for all top-$k$ norms. The reduction from any monotone symmetric norm to top-$k$ norms is deferred to Appendix A. Combining the results leads to a simultaneous $O(1)$-approximation algorithm for all monotone symmetric norms. Then in Section 4 and 5, we show how we can run the algorithm in the MPC model with nearly linear work. In particular, Section 4 and 5 discuss how to solve the LP and round the LP solution in the MPC model, respectively. The constant round MPC algorithm is described in Section 6. Theorem 1.2, 1.3, 1.4 and 1.5 are proved in Section 3, 5, 5 and 6 respectively.
@@ 3 Simultaneous $O(1)$-Approximation for Top-$k$ Norms
In the correspondent integer program, $x_{uv}$ for every $u,v\in V$ indicates if $uv$ is separated or not. We view $uv$ as an unordered pair and thus $x_{uv}$ and $x_{vu}$ are the same variable. So $(x_{uv})_{u,v\in V}$ is a metric with distances in $\{0,1\}$, which is relaxed to $[0,1]$ in the linear program. This is captured by constraints (2), (3) and (4). Notice that $\mathrm{cost}_{x}(u)$ for any $u\in V$ is a linear function of the $x$ variables. The top-$k$ norm of the fractional clustering is defined by $\mathrm{cost}^{k}_{x}=\max_{S\subseteq V:|S|=k}\sum_{u\in S}\mathrm{cost}_{x}(u)$. This could be captured by introducing a variable $z$ and constraints $z\geq\sum_{u\in S}\mathrm{cost}_{x}(u)$ for any $S\subseteq V$ of size $k$, and setting $z$ to be the objective to minimize. For simplicity, we use the form as described. Despite having exponential number of constraints, the LP can be solved efficiently as there is a simple separation oracle. Moreover, we use a combinatorial algorithm to construct a solution $x$, and thus the algorithm does not solve the LP.
@@ 3.1 Algorithm
Once we obtain Lemma 3.1, [KMZ19] provide a rounding algorithm for any feasible solution. Assume the final clustering is $\mathcal{C}_{\text{KMZ}}$. [KMZ19] ensure that for any node $u$, we have $\mathrm{cost}_{\mathcal{C}_{\text{KMZ}}}(u)\leq 5\mathrm{cost}_{x}(u)$, meaning the disagreement for $\mathcal{C}_{\text{KMZ}}$ is at most 5 times the disagreement in the LP solution. We can apply the KMZ rounding algorithm to $x_{uv}$ as output by Lemma 3.1. For any integer $k\in[n]$, let $U^{\prime}$ be the set of $k$ vertices with the largest disagreement values with respect to $\mathcal{C}_{\text{KMZ}}$. Then we have:
@@ 4.1 Line 2: Construction of Subgraph $H$
In [CALM+21], it is required that the algorithm outputs “Yes” for $uv\in E$ if $|N(u)\Delta N(v)|\leq 0.8\beta M_{uv}$ and outputs “No” for $uv\in E$ if $|N(u)\Delta N(v)|\geq\beta M_{uv}$. However, we can replace 0.8 with any constant less than 1. Additionally, the algorithm succeeds with probability $1-\frac{1}{n}$, but we can increase this to $1-\frac{1}{n^{c}}$ for any constant $c$. For completeness, we prove Theorem 4.2 in Appendix B. The theorem leads to the folowing corollary:
@@ 5.1 Rounding Algorithm
Assume that we are given a instance graph $G=(V,E)$ and a LP solution $(x_{uv})_{u,v\in V}$ such that $x_{uv}+x_{uw}+\epsilon\geq x_{vw}$ for any $u,v,w\in V^{3}$. Given a subgraph $V_{t}$ and node $w$, radius $r$, define the ball centering at $w$ with radius $r$ as $\mathrm{Ball}_{V_{t}}(w,r)=\{u\mid u\in V_{t},x_{wu}\leq r\}$. Define
@@
Our algorithm works as follows, Each round we choose a set of cluster centers and choose the ball with radius $\frac{2}{5}$ as a cluster. More precisely. At step $t$, $V_{t}$ is the set of unclustered vertices. We first compute $L_{t}^{\textmd{max}}$ to ensure that for each $u\in V_{t}$, we have $L_{t}(u)<(1+\epsilon)L_{t}^{\textmd{max}}$. For any node $u$, if $L_{t}(u)\geq L_{t}^{\textmd{max}}$, we will add $u$ to the set of candidate cluster centers $M_{t}$ (Line 6-10). Then, we compute cluster centers by adding vertices in the $M_{t}$ set to the $S_{t}$ set with probability $p_{t}$, where the more vertices in $M_{t}$ are in $\mathrm{Ball}(radius=\frac{2}{5})$ with each other, the smaller the probability (Line 11-14). After that, to avoid conflicts, we remove some cluster centers in $S_{t}$ if they are too close to each other, and derive the final cluster center set $H_{t}$ (Line 15). Let $F_{t}=\mathrm{Ball}_{V_{t}}(H_{t},\frac{2}{5})$ be the nodes clustered at step $t$. Then we add each $u\in F_{t}$ to the cluster from $H_{t}$ with minimum ID. We will remove the clustered nodes and repeat the above process until all vertices are clustered.
@@ 5.2 Approximate Ratio
Let $w$ be the first cluster center such that $x_{uw}\leq\frac{4}{5}$ when we do the above process. If all nodes in $\mathrm{Ball}_{V_{t}}(w,r)$ are still unclustered, [KMZ19] already show that the cost for short $+$edges of $u$ can be covered by the profit of $\mathrm{Ball}_{V_{t}}(w,r)$. The difficulty of the proof mainly comes from the fact that some nodes from $\mathrm{Ball}_{V_{t}}(w,r)$ may have been clustered. We will divide $x_{uw}$ into several cases and show that in all cases, we have $\mathrm{profit}_{t}(u)\geq 0$.
@@
We claim that when we consider $w$ in the above clustering process, $v$ is still unclustered. That is because $w$ is the first cluster center such that $x_{uw}\leq\frac{4}{5}$, for any cluster center $\hat{w}$ that considered before $w$, we have $x_{\hat{w}v}\geq x_{u\hat{w}}-x_{uv}-\epsilon\geq\frac{4}{5}-r-\epsilon>\frac{2}{5}$ by approximate triangle inequality. At the same time, $\hat{w}$ only includes the $\mathrm{Ball}_{V_{t}}(\hat{w},\frac{2}{5})$ into its cluster, so $\hat{w}$ will not include $v$ to its cluster. Futhermore, we claim that $w$ will include $v$ into its cluster, that’s because $x_{vw}\leq x_{uw}+x_{uv}+\epsilon\leq\frac{1}{5}+r+\epsilon\leq\frac{2}{5}$.
On the other hand, $w$ also includes $u$ into its cluster since all previous cluster center is at least $\frac{4}{5}$ away from $u$ and $x_{uw}\leq\frac{1}{5}$. Combining them together, we know that $uv$ will be in the same cluster and $ALG(uv)=0$, which implies $\mathrm{profit}_{t}(uv)\geq x_{uv}\geq 0$. ∎
By Lemma 5.3 and 5.4, we know that when $w$ does not exists or $x_{uw}\leq\frac{1}{5}$, we have $\mathrm{profit}_{t}(u)=\sum_{uv\in\Delta E_{t}}\mathrm{profit}_{t}(uv)\geq 0$. For the remaining cases, we will follow the blue-print of the proof in [KMZ19]. Note that $\mathrm{profit}_{t}(uv)<0$ only if $uv\in E$ and $x_{uv}\leq r$. We will charge the negative profit to the profit of $\mathrm{Ball}_{V_{t}}(w,r)$. More precisely, let
@@
Since $v\in\mathrm{Ball}_{V_{t}}(w,\frac{2}{5})$, $v$ must be clustered after we consider $w$, while $u$ is still not clustered at this time. Therefore, we know that $uv$ must be in different cluster and $ALG(uv)=0$ and
@@
Combining the profit for $P_{\textrm{high}}(u)$ and $P_{\textrm{low}}(u)$, we can deduct that the profit for any node $u$ is non-negative when $x_{uw}\in(\frac{1}{5},\frac{4}{5}]$
@@ 5.3 Running time
At some round $s$, let $L^{\max}_{s}$ and $\Delta_{s}$ be the value we set up at lines 7 and 11. Then, after $O(\log n)$ rounds, at round $e=s+O(\log n)$, with high probability, either $L^{\max}_{e}<L^{\max}_{s}$ or $\Delta_{e}\leq\frac{\Delta_{s}}{2}$.
@@ 6.2 Approximate Ratio
Recall that ${\mathcal{C}}$ and ${\mathcal{C}}^{*}$ represent the top-$k$ optimal clustering for graph $G$ and graph $G_{2}$, respectively. Let ${\mathrm{opt}}^{k}$ denote the cost of the optimum solution under the top-$k$ norm for $G$. When we transfer the graph $G_{2}$ to $G$, we will incur an increase in the cost of $F$. Lemma 6.4 provide the upper bounds on the increase in the cost of $F$, given by:

### v2

@@ 1 Introduction
In light of the growing networks, it is imperative to develop efficient parallel algorithms. This urgency is particularly pronounced in machine learning and data mining applications, where timely and efficient processing is essential for extracting meaningful insights from vast datasets. Many works in the literature aim to design efficient parallel algorithms [BFS12, CDK14, PPO+15, FN19, CCMU21, CALM+21, AW22, CKL+24, CHS24]. The MPC model, as a theoretical abstraction of several real-world parallel models such as MapReduce [DG08], is a prevalent methodology employed in these works.
@@ 1.2 Overview of Techniques
By [CS19], we can reduce the problem of approximating all monotone symmetric norms to approximating all top-$k$ norms. We then construct a fractional solution $x$, which is a metric over $V$ with range $[0,1]$, such that the fractional disagreement vector for $x$ has top-$k$ norm at most $12.66\cdot{\mathrm{opt}}_{k}$ for any $k\in[n]$, where ${\mathrm{opt}}_{k}$ is the cost of the optimum clustering under the top-$k$ norm. Then, we can use the 5-approximate rounding algorithm of KMZ [KMZ19], to obtain a simultaneous $63.3$-approximation for all top-$k$ norms. The KMZ rounding algorithm has two crucial properties that we need: it does not depend on $k$ and it achieves a per-vertex guarantee.
We elaborate more on how to construct the metric $x:{V\choose 2}\to[0,1]$. A natural idea to assign the LP values, that was used by [DMN24], is to set $x_{uv}$ based on the intersection of the neighborhood between $u$ and $v$. Intuitively, the more common neighbors two nodes share, the closer they should be. A straightforward approach to implementing this idea is to set $x_{uv}=1-\frac{|N(u)\cap N(v)|}{\max(d(u),d(v))}$, where $N(u)$ denotes the neighboring nodes of $u$ in $G$ and $d(u)=|N(u)|$ denotes the degree of $u$; it is convenient to assume $u\in N(u)$. This approach works for the top-$1$ norm (i.e., the $\ell_{\infty}$ norm) as discussed in [DMN23], but fails for the top-$n$ norm (i.e., the $\ell_{1}$ norm). Consider a star graph, where the optimal clustering under the top-$n$ norm has a cost of $n-2$. This approach will assign $x_{uv}=1-\frac{1}{2}=1/2$ for all --edges, leading to an LP cost of $\Theta(n^{2})$ and a gap of $\Omega(n)$. [DMN24] addressed the issue by rounding up LP values to $1$ for --edge, if for a given node, its total --edges LP disagreement is larger than the number of its $+$edges. After the transformation, the triangle inequalities are only satisfied approximately, but this can be handled with $O(1)$ loss in the approximation ratio.
@@
The challenge to implement the algorithm in MPC model is the sequential nature of the algorithm. [KMZ19] observes that in each round, if we select the nodes that maximize $L(u)=\sum_{x_{uv}\leq r}(r-x_{uv})$ as cluster center, we can effectively bound each node’s algorithmic cost, where $r$ is the final ratio. However, choosing a node that maximizes some target inherently makes the process sequential. Our key observation is that, instead of selecting the node that maximizes $L(u)$, we can allow some approximation. This strategy still permits achieving a reasonable approximate ratio with an additional $1+\epsilon$ overhead while allowing the selection of multiple nodes as cluster centers, thereby parallelizing the rounding process. In each round, there might be several candidate cluster centers with conflicts. To resolve these conflicts, we employ the classical Luby’s algorithm [Lub85, CDK14] to find a maximal independent set, ensuring that none of the cluster centers have conflicts.
@@
We give some preliminary remarks in Section 2. In Section 3, we describe our simultaneous $O(1)$-approximation algorithm for correlation clustering for all top-$k$ norms. The reduction from any monotone symmetric norm to top-$k$ norms is deferred to Appendix A. Combining the results leads to a simultaneous $O(1)$-approximation algorithm for all monotone symmetric norms. Then in Section 4 and 5, we show how we can run the algorithm in the MPC model with nearly linear work. In particular, Section 4 and 5 discuss how to solve the LP and round the LP solution in the MPC model, respectively. The constant round MPC algorithm is described in Section 6. Theorem 1.2, 1.3, 1.4 and 1.5 are proved in Section 3, 5, 5 and 6 respectively.
@@ 3 Simultaneous $O(1)$-Approximation for Top-$k$ Norms
In the correspondent integer program, $x_{uv}$ for every $u,v\in V$ indicates if $uv$ is separated or not. We view $uv$ as an unordered pair and thus $x_{uv}$ and $x_{vu}$ are the same variable. So $(x_{uv})_{u,v\in V}$ is a metric with distances in $\{0,1\}$, which is relaxed to $[0,1]$ in the linear program. This is captured by constraints (2), (3) and (4). Notice that $\mathrm{cost}_{x}(u)$ for any $u\in V$ is a linear function of the $x$ variables. The top-$k$ norm of the fractional clustering is defined by $\mathrm{cost}^{k}_{x}=\max_{S\subseteq V:|S|=k}\sum_{u\in S}\mathrm{cost}_{x}(u)$. This could be captured by introducing a variable $z$ and constraints $z\geq\sum_{u\in S}\mathrm{cost}_{x}(u)$ for any $S\subseteq V$ of size $k$, and setting $z$ to be the objective to minimize. For simplicity, we use the form as described. Despite having an exponential number of constraints, the LP can be solved efficiently as there is a simple separation oracle. Moreover, we use a combinatorial algorithm to construct a solution $x$, and thus the algorithm does not solve the LP.
@@ 3.1 Algorithm
Once we obtain Lemma 3.1, [KMZ19] provides a rounding algorithm for any feasible solution. Assume the final clustering is $\mathcal{C}_{\text{KMZ}}$. [KMZ19] ensures that for any node $u$, we have $\mathrm{cost}_{\mathcal{C}_{\text{KMZ}}}(u)\leq 5\mathrm{cost}_{x}(u)$, meaning the disagreement for $\mathcal{C}_{\text{KMZ}}$ is at most 5 times the disagreement in the LP solution. We can apply the KMZ rounding algorithm to $x_{uv}$ as output by Lemma 3.1. For any integer $k\in[n]$, let $U^{\prime}$ be the set of $k$ vertices with the largest disagreement values with respect to $\mathcal{C}_{\text{KMZ}}$. Then we have:
@@ 4.1 Line 2: Construction of Subgraph $H$
In [CALM+21], it is required that the algorithm outputs “Yes” for $uv\in E$ if $|N(u)\Delta N(v)|\leq 0.8\beta M_{uv}$ and outputs “No” for $uv\in E$ if $|N(u)\Delta N(v)|\geq\beta M_{uv}$. However, we can replace 0.8 with any constant less than 1. Additionally, the algorithm succeeds with probability $1-\frac{1}{n}$, but we can increase this to $1-\frac{1}{n^{c}}$ for any constant $c$. For completeness, we prove Theorem 4.2 in Appendix B. The theorem leads to the following corollary:
@@ 5.1 Rounding Algorithm
Assume that we are given an instance graph $G=(V,E)$ and a LP solution $(x_{uv})_{u,v\in V}$ such that $x_{uv}+x_{uw}+\epsilon\geq x_{vw}$ for any $u,v,w\in V^{3}$. Given a subgraph $V_{t}$ and node $w$, radius $r$, define the ball centering at $w$ with radius $r$ as $\mathrm{Ball}_{V_{t}}(w,r)=\{u\mid u\in V_{t},x_{wu}\leq r\}$. Define
@@
Our algorithm works as follows: in each round, we choose a set of cluster centers and choose the ball with radius $\frac{2}{5}$ as a cluster. More precisely. At step $t$, $V_{t}$ is the set of unclustered vertices. We first compute $L_{t}^{\textmd{max}}$ to ensure that for each $u\in V_{t}$, we have $L_{t}(u)<(1+\epsilon)L_{t}^{\textmd{max}}$. For any node $u$, if $L_{t}(u)\geq L_{t}^{\textmd{max}}$, we will add $u$ to the set of candidate cluster centers $M_{t}$ (Line 6-10). Then, we compute cluster centers by adding vertices in the $M_{t}$ set to the $S_{t}$ set with probability $p_{t}$, where the more vertices in $M_{t}$ are in $\mathrm{Ball}(radius=\frac{2}{5})$ with each other, the smaller the probability (Line 11-14). After that, to avoid conflicts, we remove some cluster centers in $S_{t}$ if they are too close to each other, and derive the final cluster center set $H_{t}$ (Line 15). Let $F_{t}=\mathrm{Ball}_{V_{t}}(H_{t},\frac{2}{5})$ be the nodes clustered at step $t$. Then we add each $u\in F_{t}$ to the cluster from $H_{t}$ with minimum ID. We will remove the clustered nodes and repeat the above process until all vertices are clustered.
@@ 5.2 Approximate Ratio
Let $w$ be the first cluster center such that $x_{uw}\leq\frac{4}{5}$ when we do the above process. If all nodes in $\mathrm{Ball}_{V_{t}}(w,r)$ are still unclustered, [KMZ19] already shows that the cost for short $+$edges of $u$ can be covered by the profit of $\mathrm{Ball}_{V_{t}}(w,r)$. The difficulty of the proof mainly comes from the fact that some nodes from $\mathrm{Ball}_{V_{t}}(w,r)$ may have been clustered. We will divide $x_{uw}$ into several cases and show that in all cases, we have $\mathrm{profit}_{t}(u)\geq 0$.
@@
We claim that when we consider $w$ in the above clustering process, $v$ is still unclustered. That is because $w$ is the first cluster center such that $x_{uw}\leq\frac{4}{5}$, for any cluster center $\hat{w}$ that considered before $w$, we have $x_{\hat{w}v}\geq x_{u\hat{w}}-x_{uv}-\epsilon\geq\frac{4}{5}-r-\epsilon>\frac{2}{5}$ by approximate triangle inequality. At the same time, $\hat{w}$ only includes the $\mathrm{Ball}_{V_{t}}(\hat{w},\frac{2}{5})$ into its cluster, so $\hat{w}$ will not include $v$ to its cluster. Furthermore, we claim that $w$ will include $v$ into its cluster, that’s because $x_{vw}\leq x_{uw}+x_{uv}+\epsilon\leq\frac{1}{5}+r+\epsilon\leq\frac{2}{5}$.
On the other hand, $w$ also includes $u$ into its cluster since all previous cluster centers are at least $\frac{4}{5}$ away from $u$ and $x_{uw}\leq\frac{1}{5}$. Combining them together, we know that $uv$ will be in the same cluster and $ALG(uv)=0$, which implies $\mathrm{profit}_{t}(uv)\geq x_{uv}\geq 0$. ∎
By Lemma 5.3 and 5.4, we know that when $w$ does not exist or $x_{uw}\leq\frac{1}{5}$, we have $\mathrm{profit}_{t}(u)=\sum_{uv\in\Delta E_{t}}\mathrm{profit}_{t}(uv)\geq 0$. For the remaining cases, we will follow the blue-print of the proof in [KMZ19]. Note that $\mathrm{profit}_{t}(uv)<0$ only if $uv\in E$ and $x_{uv}\leq r$. We will charge the negative profit to the profit of $\mathrm{Ball}_{V_{t}}(w,r)$. More precisely, let
@@
Since $v\in\mathrm{Ball}_{V_{t}}(w,\frac{2}{5})$, $v$ must be clustered after we consider $w$, while $u$ is still not clustered at this time. Therefore, we know that $uv$ must be in different clusters and $ALG(uv)=0$ and
@@
Combining the profit for $P_{\textrm{high}}(u)$ and $P_{\textrm{low}}(u)$, we can deduce that the profit for any node $u$ is non-negative when $x_{uw}\in(\frac{1}{5},\frac{4}{5}]$
@@ 5.3 Running time
At some round $s$, let $L^{\max}_{s}$ and $\Delta_{s}$ be the values we set up at lines 7 and 11. Then, after $O(\log n)$ rounds, at round $e=s+O(\log n)$, with high probability, either $L^{\max}_{e}<L^{\max}_{s}$ or $\Delta_{e}\leq\frac{\Delta_{s}}{2}$.
@@ 6.2 Approximate Ratio
Recall that ${\mathcal{C}}$ and ${\mathcal{C}}^{*}$ represent the top-$k$ optimal clustering for graph $G$ and graph $G_{2}$, respectively. Let ${\mathrm{opt}}^{k}$ denote the cost of the optimum solution under the top-$k$ norm for $G$. When we transfer the graph $G_{2}$ to $G$, we will incur an increase in the cost of $F$. Lemma 6.4 provides the upper bounds on the increase in the cost of $F$, given by: